OpenStudy (anonymous):

Put the equation in the form ax2 + bx + c = 0. Enter exponents using the caret ( ^ ). For example, you would enter 4x2 as 4x^2. x2 + 6x = - 4x2 + 2

OpenStudy (texaschic101):

x^2 + 6x = -4x^2 + 2 (add 4x^2 to both sides) x^2 + 4x^2 + 6x = 2 (now subtract 2 from both sides) 5x^2 + 6x - 2 = 0 <-- answer